You’re considering investing $35,000 in a new development project which you expect to complete in five...

  1. You’re considering investing $35,000 in a new development project which you expect to complete in five years. You’ll earn all your return when you sell the project (i.e., you earn nothing in years 1-4). Answer the following:
    1. You’ve done some research and similar projects with similar risks usually earn 16% annual interest. Assuming that you have the option to invest in one of the other projects instead, what is the minimum sales price you would need to receive in year five for you to accept this project?
Answer #1

minimum sales price would be future value of the project at 16% annual interest for 5 years.

minimum sales price = investment value*(1+annual interest rate)life of project = $35,000*(1+0.16)5 = $35,000*1.165 = $35,000*2.1003416576‬ = $73,511.96 or $73,512